OpenStudy (anonymous):

If x^2 + 2xy + y^2 = 25, x + y > 0 and x - y = 1, then x =

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

x^2 + 2xy + y^2=(x+y)^2 So, how can we use that?

OpenStudy (anonymous):

i guess you can solve for y and then plug in

OpenStudy (anonymous):

I had this idea: x^2 + 2xy + y^2=25 So then: (x+y)^2=25 So: x+y=5 OR x+y=-5 -5 can't be because x+y>0 So then: x+y=5 And using this: x - y = 1 We add them: 2x=6 X=3 Makes sense?

OpenStudy (anonymous):

@lawls Does this makes sense? Feel free to ask! :)

OpenStudy (anonymous):

yeah it does. i got the same answer. thanks :)

OpenStudy (anonymous):

You are welcome! Glad to help you! :)